$165000 in tickets were sold, 8000 tickets were sold. How many for $15 and how many for $25 were sold?
Answer by checkley77(12844)

25X+15(8,000-X)=165,000
25X+120,000-15X=165,000
10X=165,000-120,000
10X=45,000
X=45,000/10
X=4,500 $25 TICKETS WERE SOLD.
8,000-4,500=3,500 $15 TICKETS SOLD.
PROOF:
25*4,500+15*3,500=165,000
112,500+52,500=165,000
165,000=165,000